OpenStudy (anonymous):

A smokestack deposits soot on the ground with a concentration inversely proportional to the square of the distance from the stack. With two smokestacks d miles apart, the concentration of the combined deposits on the line joining them, at a distance x from one stack, is given by S = cx^-2+k(d-x)^2 k=9c find minimum need help with back half of question, please

OpenStudy (anonymous):

What is c?

OpenStudy (anonymous):

where c and k are positive constants which depend on the quantity of smoke each stack is emitting. If k = 9 c, find the point on the line joining the stacks where the concentration of the deposit is a minimum.

OpenStudy (anonymous):

So I get S= cx^-2+9c(d-x)^-2 and factor out the c for later to get x^-2+9(d-x)^-2

OpenStudy (anonymous):

Well, first you need to take the derivative with respect to x. Once you have that function, set it equal to 0 and solve for x.

OpenStudy (anonymous):

yup. I get x=d/3.080083823, right?

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

Where did c go?

OpenStudy (anonymous):

so, cd/3.080083823

OpenStudy (anonymous):

all these variables really screw me up

OpenStudy (anonymous):

ok, what do you get as the derivative of S(x)

OpenStudy (anonymous):

-2x^-3+18(d-x)^-3

OpenStudy (anonymous):

ok you're close. For the power rule of differentiation, you multiply by the power, then lower the exponent.

OpenStudy (anonymous):

where didnt I do that? I even applied the chain rule for the second

OpenStudy (anonymous):

oops, I read it wrong. Sorry about that. What you have is right.

OpenStudy (anonymous):

so what do u get for x? and what do I do with that info?

OpenStudy (anonymous):

So setting that equal to 0, you get 2c/(x^3) = 18c/(d-x)^3 1/x^3 = 9/(d-x)^3 (d-x)^3/x^3 = 9 Solve that for x and your answer should only be in terms of d. Sorry. I think I should be the one asking the question =p

OpenStudy (anonymous):

Take the x value you get that is in terms of d, and plug it back into the original equation to find the minimum value for S(x).

OpenStudy (anonymous):

Well the deadline was midnight and I had it right the whole time just entered it wrong.. :(

OpenStudy (anonymous):

d/(1+2.08)

OpenStudy (anonymous):

thanks for the feedback, seriously!

OpenStudy (anonymous):

Haha, glad I could "help".